OpenStudy (anonymous):

If 7x + 3y = 10 then what is the value of 3x?

OpenStudy (ikram002p):

7x=10-3y thus 3x=10-3y-4x ? you want it like this ?

OpenStudy (eric_d):

 Maybe, our nxt step is... substituting the above equation into \(3x\)